# A Backpropagation-Free Feedback–Hebbian Network for Continual Learning Dynamics

### v1

@@ III-B Sequential AB: Retention and LTD-Like Unlearning
We next test unlearning and retention under the sequential AB protocol (Regime I; Sec. II-D). In the trajectory plots, the Phase A/Phase B boundary and the baseline and endpoint epochs are marked on the highlighted traces.
LTD-like unlearning in the forward output stage.
Fig. 4 plots the output connectivity of forward layer 2. During Phase A, connectivity concentrates on the pair A target outputs (sites 8 and 9). After switching to Phase B, connectivity to the pair B targets (sites 5 and 6) increases, while the earlier pair A targets are driven down, consistent with LTD-like suppression at the output stage. This is summarized by the retention index (Eq. (2)), where indicates minimal Phase B change and the sign indicates increase/decrease: at the pair B targets, and , whereas at the pair A targets, and .
Retention of the earlier association in the feedback pathway.
Fig. 5 plots the input connectivity of feedback layer 2 over the same training sequence. In contrast to the forward output layer, the pair A-related sites remain stable during Phase B: and , indicating minimal Phase B change relative to the Phase A change.
Sequential training therefore yields a separation of roles across pathways: the forward output stage shows an LTD-like reduction of the earlier association during Phase B, while the feedback pathway preserves an A-related trace during acquisition of the new association.
@@ III-C Interleaved A/B: Conditioning Under Temporal Alternation
We next evaluate the conditioning primitive under temporal alternation (Regime II; Sec. II-D), where samples from pair A and pair B are presented in a deterministic interleaving within each epoch. In this setting, conditioning corresponds to concurrent acquisition and maintenance of both associations rather than sequential suppression of the earlier mapping.
Concurrent co-maintenance at the forward output stage.
Fig. 6 plots the output connectivity of forward layer 2 under interleaved training. Connectivity to the pair A target outputs (sites 8 and 9) and to the pair B target outputs (sites 5 and 6) increases during training and remains elevated relative to non-target sites. Unlike the sequential AB protocol (Sec. III-B), the interleaved schedule does not induce a sustained depression of the earlier association; instead, both target sets are jointly maintained at the output stage.
Parallel co-learning in the feedback pathway.
@@
Together, these results show that temporal alternation supports concurrent encoding of both associations in the forward prediction and feedback pathways, consistent with the operational definition of conditioning introduced above.
@@ III-D Architectural Controls: Depth and the Role of Feedback
We next evaluate architectural controls to assess whether the continual-learning–relevant primitives reported above depend on network depth and on the presence of a dedicated feedback pathway. Fig. 8 summarizes four architectures across three probes.
@@
The Forward prediction and Regeneration columns are evaluated after the sequential AB protocol (Regime I; Sec. II-D) at the end of Phase B, so correct behavior is characterized by a selective expression of the pair B mapping at the output and, when feedback is present, regeneration consistent with the pair B association. The Conditioning column is evaluated after the interleaved A/B protocol (Regime II), where correct behavior requires concurrent maintenance of both pair A and pair B target outputs.
Shallow architecture (1F+1B) is insufficient.
@@
In the 1F+1B row of Fig. 8, the network does not consistently exhibit the selective patterns required by the three probes: post Phase B forward prediction is not cleanly dominated by the pair B targets, regeneration is not sharply aligned with the corresponding association, and the interleaved conditioning probe fails to maintain both mappings concurrently. This indicates that a single forward/feedback stage is insufficient to support the separation of roles required for stable supervised prediction together with feedback-supported reconstruction in this setting.
Proposed architecture (2F+2B) supports all probes.
@@
In contrast, the proposed 2F+2B architecture exhibits the expected behaviors across all columns in Fig. 8. After sequential training, forward prediction at the end of Phase B emphasizes the pair B target outputs (sites 5 and 6) while suppressing non-target sites, and the regeneration probe yields a selective reconstructed pattern consistent with the pair B association. Under interleaved training, the conditioning probe shows concurrent expression of both pair A targets (sites 8 and 9) and pair B targets (sites 5 and 6), consistent with co-maintenance under temporal alternation.
Deeper architecture (3F+3B) is not necessary on this task.
@@
The 3F+3B architecture produces qualitatively similar outcomes to 2F+2B across the same three probes (Fig. 8). For the present two-pair dataset, additional depth does not yield an observable change in the forward prediction, regeneration, or conditioning behaviors, supporting the use of 2F+2B as a minimal sufficient configuration.
Removing feedback (2F-only) breaks regeneration and conditioning.
@@
Finally, the 2F-only architecture isolates the role of feedback. As expected, the regeneration probe fails in the absence of a feedback pathway. More importantly, the conditioning probe no longer maintains both associations: after interleaved training, one target set dominates while the other collapses toward non-target levels (Fig. 8, 2F-only row). Notably, post Phase B forward prediction can still emphasize the pair B targets, indicating that the forward pathway alone can express the most recently trained association, whereas dedicated feedback is required for regeneration and for stabilizing concurrent associations under temporal alternation.
@@
Together, these architectural controls indicate that two forward and two dedicated feedback layers are sufficient and, for the present task, effectively minimal: shallower networks fail to exhibit the full set of primitives, deeper networks do not add qualitative benefits, and removing feedback selectively abolishes regeneration and conditioning.
@@ III-E Rule-Term Ablations: Covariance, Oja Normalization, and Supervision
We next ablate individual terms in the local update rule (Eq. (1)) to assess which components are necessary for the continual-learning–relevant primitives observed under the full rule. As references for the desired behaviors, sequential AB training under the proposed update exhibits LTD-like suppression of the earlier association at the forward output stage with retention in feedback (Figs. 4–5), and interleaved training maintains both associations concurrently (Figs. 6–7).
@@
Weight-scale effects of normalization and centering.
Table I reports the maximum absolute weight across all learned matrices at the end of sequential AB training (end of Phase B) for four update variants. Removing the Oja-style decay term increases the peak weight magnitude relative to the full rule (1.72 vs. 1.16), and removing the covariance (centering) term yields the largest peak magnitude (1.99). In our experiments, these larger weight scales coincide with slower suppression of earlier-association connectivity during Phase B (as reflected by the forward output connectivity trajectories), and are consistent with the common view that larger synaptic gains can make local learning dynamics harder to reverse and potentially less stable, particularly in feedback-driven settings.
Removing the supervised term disrupts unlearning under sequential training.
We next remove the local supervised drive from Eq. (1) while keeping the covariance and Oja terms unchanged. Under sequential AB training, the forward layer 2 output connectivity no longer shows a clean LTD-like reduction of the earlier association during Phase B (Fig. 9). In contrast to the full rule (Fig. 4), the earlier pair A target outputs (sites 8–9) are not effectively driven down while pair B targets (sites 5–6) are acquired. This indicates that Hebbian covariance updates with normalization are not sufficient on their own to suppress non-target outputs at the output stage; the supervised term provides the local error-driven component needed to selectively depress activity inconsistent with the current targets.
Removing the supervised term also impairs conditioning under interleaving.
The same ablation also disrupts conditioning under temporal alternation. Under the interleaved A/B protocol (Regime II), forward layer 2 output connectivity fails to maintain both target sets concurrently when the supervised term is removed (Fig. 10). Whereas the full rule preserves elevated connectivity at both pair A targets (sites 8–9) and pair B targets (sites 5–6) throughout training (Fig. 6), removing supervision leads to dominance or interference that prevents stable co-maintenance of the two associations. Together with the sequential results, this supports the functional role of the supervised drive as the mechanism that enforces output selectivity—enabling LTD-like unlearning of outdated associations and supporting concurrent conditioning under alternation.
@@ IV Conclusion
Across two training regimes, the model exhibited three operational continual-learning–relevant primitives. First, under sequential training, forward layer 2 output connectivity shifted from the pair A targets (sites 8–9) to the pair B targets (sites 5–6), while the earlier pair A structure was selectively suppressed, consistent with an LTD-like unlearning effect at the output stage. Second, in the same sequential protocol, feedback layer 2 preserved an A-related trace: input connectivity at the pair A-related sites remained near-stable during Phase B, indicating retention despite the forward pathway’s adaptation to the newly trained mapping. Third, under deterministic interleaving , both associations were concurrently maintained at the forward output stage, with parallel strengthening and persistence in the feedback pathway, consistent with conditioning under temporal alternation rather than sequential suppression.
@@
Architectural controls and rule-term ablations further sharpened the mechanistic interpretation. A single forward/feedback stage was insufficient to support the full set of probes, whereas the proposed 2F+2B configuration reliably expressed forward prediction, feedback-driven regeneration, sequential unlearning with retention, and interleaved co-maintenance. Removing the feedback pathway abolished regeneration and degraded stable co-maintenance under alternation, isolating dedicated feedback as the component that enables regeneration and supports concurrent associations. At the learning-rule level, covariance and Oja-style stabilization regulated synaptic scale and supported stable adaptation, while removing the local supervised drive disrupted output selectivity, impairing both LTD-like unlearning under sequential training and co-maintenance under interleaving. Together, these findings support a division of labor in which forward layers express the currently trained output mapping, while feedback layers can preserve and propagate association-related traces that are not forced to mirror the immediate output targets.
@@
This study is intentionally minimal, and several limitations motivate future work. The present task involves two low-dimensional associations and short training horizons; it remains to be tested how the same local learning dynamics behave under higher-dimensional inputs, longer task sequences, structured task overlap, and noisier regimes. A natural next step is to extend the same controlled sequential and interleaved protocols to richer multi-class or multi-label datasets while retaining diagnostics that localize where retention and interference emerge across layers and pathways. More broadly, dedicated feedback trained by local rules offers a flexible design space for exploring additional biologically motivated motifs (e.g., alternative reconstruction targets, partial feedback, sparsity or gating) without introducing global gradient transport. Within this scope, an important open question is when such local, feedback-trained architectures can serve as a useful complement to gradient-based training in settings where locality, interpretability of dynamics, or circuit-level constraints are central.

### v2

@@ III-B Sequential AB: Retention and LTD-Like Unlearning
Next, we test unlearning and retention under the sequential AB protocol (Regime I). In the trajectory plots, the Phase 1/Phase 2 boundary and the baseline and endpoint epochs are marked on the highlighted traces.
LTD-like unlearning in the forward output stage
Fig. 4 plots the output connectivity of forward layer 2. During Phase 1, connectivity concentrates on the pair A target outputs (sites 8 and 9). After switching to Phase 2, connectivity to the pair B targets (sites 5 and 6) increases, while the earlier pair A targets are driven down, consistent with LTD-like suppression at the output stage. This is summarized by the retention index (Eq. (2)), where indicates minimal Phase 2 change and the sign indicates increase/decrease: at the pair B targets, and , whereas at the pair A targets, and .
Retention of the earlier association in the feedback pathway
Fig. 5 plots the input connectivity of feedback layer 2 over the same training sequence. In contrast to the forward output layer, the pair A-related sites remain stable during Phase 2: and , indicating minimal Phase 2 change relative to the Phase 1 change.
Sequential training therefore yields a separation of roles across pathways: the forward output stage shows an LTD-like reduction of the earlier association during Phase 2, while the feedback pathway preserves an A-related trace during acquisition of the new association.
@@ III-C Interleaved A/B: Conditioning Under Temporal Alternation
We then evaluate the conditioning primitive under temporal alternation (Regime II), where samples from pair A and pair B are presented in a deterministic interleaving within each epoch. In this setting, conditioning corresponds to concurrent acquisition and maintenance of both associations rather than sequential suppression of the earlier mapping.
Concurrent co-maintenance at the forward output stage
Fig. 6 plots the output connectivity of forward layer 2 under interleaved training. Connectivity to the pair A target outputs (sites 8 and 9) and to the pair B target outputs (sites 5 and 6) increases during training and remains elevated relative to non-target sites. Unlike the sequential AB protocol analyzed above, the interleaved schedule does not induce a sustained depression of the earlier association; instead, both target sets are jointly maintained at the output stage.
Parallel co-learning in the feedback pathway
@@
Together, these results show that temporal alternation supports concurrent encoding of both associations in the forward prediction and feedback pathways.
@@ III-D Architectural Controls: Depth and the Role of Feedback
Afterward, we evaluate architectural controls to assess whether the continual-learning–relevant primitives reported above depend on network depth and on the presence of a dedicated feedback pathway. Fig. 8 summarizes four architectures across three probes.
@@
The Forward prediction and Regeneration columns are evaluated after the sequential AB protocol at the end of Phase 2, so correct behavior is characterized by a selective expression of the pair B mapping at the output and, when feedback is present, regeneration consistent with the pair B association. The Conditioning column is evaluated after the interleaved A/B protocol, where correct behavior requires concurrent maintenance of both pair A and pair B target outputs.
Shallow architecture (1FF+1FB) is insufficient
@@
In the 1FF+1FB as depicted in Fig. 8, the network does not consistently exhibit the selective patterns required by the three probes: post Phase 2 forward prediction is not cleanly dominated by the pair B targets, regeneration is not sharply aligned with the corresponding association, and the interleaved conditioning probe fails to maintain both mappings concurrently. This indicates that a single forward/feedback stage is insufficient to support the separation of roles required for stable supervised prediction together with feedback-supported reconstruction in this setting.
Proposed architecture (2FF+2FB) supports all probes
@@
In contrast, the proposed 2FF+2FB architecture exhibits the expected behaviors across all columns as shown in Fig. 8. After sequential training, forward prediction at the end of Phase 2 emphasizes the pair B target outputs (sites 5 and 6) while suppressing non-target sites, and the regeneration probe yields a selective reconstructed pattern consistent with the pair B association. Under interleaved training, the conditioning probe shows concurrent expression of both pair A targets (sites 8 and 9) and pair B targets (sites 5 and 6), consistent with co-maintenance under temporal alternation.
Deeper architecture (3FF+3FB) is not necessary on this task
@@
The 3FF+3FB architecture produces qualitatively similar outcomes to 2FF+2FB across the same three probes, as also described in Fig. 8. For the present two-pair dataset, additional depth does not yield an observable change in the forward prediction, regeneration, or conditioning behaviors, supporting the use of 2FF+2FB as a minimal sufficient configuration.
Removing feedback (2FF-only) breaks regeneration and conditioning
@@
Finally, the 2FF-only architecture isolates the role of feedback. As expected, the regeneration probe fails due to the absence of a feedback pathway. More importantly, the conditioning probe can no longer maintain both associations: after interleaved training, one target set dominates while the other collapses toward non-target levels as highlighted in Fig. 8 also. Notably, post Phase 2 forward prediction can still emphasize the pair B targets, indicating that the forward pathway alone can express the most recently trained association, whereas dedicated feedback is required for regeneration and for stabilizing concurrent associations under temporal alternation.
@@
In short, these architectural controls indicate that two forward and two dedicated feedback layers are sufficient and, for the present task, effectively minimal. Shallower networks fail to exhibit the full set of primitives while deeper networks do not add qualitative benefits. Last but not least, removing feedback selectively abolishes regeneration and conditioning.
@@ III-E Rule-Term Ablations: Covariance, Oja Normalization, and Supervision
Lastly, we ablate individual terms in the local update rule (Eq. (1)) to assess which components are necessary for the continual-learning–relevant primitives observed under the full rule. As references for the desired behaviors, sequential AB training under the proposed update exhibits LTD-like suppression of the earlier association at the forward output stage with retention in feedback as depicted in Figs. 4, 5, and interleaved training maintains both associations concurrently as highlighted in Figs. 6, 7.
@@
Weight-scale effects of normalization and centering
Table I reports the maximum absolute weight across all learned matrices at the end of sequential AB training (end of Phase 2) for four update variants. Removing the Oja-style decay term increases the peak weight magnitude relative to the full rule (1.72 vs. 1.16), and removing the covariance (centering) term yields the largest peak magnitude (1.99). In our experiments, these larger weight scales coincide with slower suppression of earlier-association connectivity during Phase 2 (as reflected by the forward output connectivity trajectories), and are consistent with the common view that larger synaptic gains can make local learning dynamics harder to reverse and potentially less stable, particularly in feedback-driven settings.
Removing the supervised term disrupts unlearning under sequential training
We then remove the local supervised drive from Eq. (1) while keeping the covariance and Oja terms unchanged. Under sequential AB training, the forward layer 2 output connectivity no longer shows a clean LTD-like reduction of the earlier association during Phase 2 as plotted in Fig. 9. In contrast to the full rule as depicted in Fig. 4, the earlier pair A target outputs (sites 8–9) are not effectively driven down while pair B targets (sites 5–6) are acquired. This indicates that Hebbian covariance updates with normalization are not sufficient on their own to suppress non-target outputs at the output stage; the supervised term provides the local error-driven component needed to selectively depress activity inconsistent with the current targets.
Removing the supervised term also impairs conditioning under interleaving
The same ablation also disrupts conditioning under temporal alternation. Under the interleaved A/B protocol (Regime II), forward layer 2 output connectivity fails to maintain both target sets concurrently when the supervised term is removed as shown in Fig. 10. Whereas the full rule preserves elevated connectivity at both pair A targets (sites 8–9) and pair B targets (sites 5–6) throughout training as compared in Fig. 6, removing supervision leads to dominance or interference that prevents stable co-maintenance of the two associations. Together with the sequential results, this supports the functional role of the supervised drive as the mechanism that enforces output selectivity—enabling LTD-like unlearning of outdated associations and supporting concurrent conditioning under alternation.
@@ IV Conclusions
Across two training regimes, the model exhibited three operational continual-learning–relevant primitives. First, under sequential training, forward layer 2 output connectivity shifted from the pair A to pair B, while the earlier pair A structure was selectively suppressed, consistent with an LTD-like unlearning effect at the output stage. Secondly, in the same sequential protocol, feedback layer 2 preserved an A-related trace: input connectivity at the pair A-related sites remained near-stable during Phase B, indicating retention despite the forward pathway’s adaptation to the newly trained mapping. Lastly, under alternating sequence, both associations were concurrently maintained at the forward output stage, with parallel strengthening and persistence in the feedback pathway, consistent with conditioning under temporal alternation rather than sequential suppression.
@@
Architectural controls and rule-term ablations further sharpened the mechanistic interpretation. A 1FF+1FB was insufficient to support the full set of probes, whereas the proposed 2FF+2BB configuration reliably expressed forward prediction, feedback-driven regeneration, sequential unlearning with retention, and interleaved co-maintenance. Removing the feedback pathway abolished regeneration and degraded stable co-maintenance under alternation, isolating dedicated feedback as the component that enables regeneration and supports concurrent associations. At the learning-rule level, covariance and Oja-style stabilization regulated synaptic scale and supported stable adaptation, while removing the local supervised drive disrupted output selectivity, impairing both LTD-like unlearning under sequential training and co-maintenance under alternating sequence. In short, these findings support a division of labor in which forward layers express the currently trained output mapping, while feedback layers can preserve and propagate association-related traces that are not forced to mirror the immediate output targets.
V Outlook
@@
This study is intentionally minimal, and several limitations motivate future work. The present task comprises two low-dimensional associations and short training horizons; it therefore remains unclear how the same local learning dynamics scale to higher-dimensional inputs, longer task sequences, structured task overlap, and noisier settings. A natural next step is to extend the same controlled sequential and interleaved protocols to richer multi-class or multi-label datasets while retaining diagnostics that localize where retention and interference emerge across layers and pathways. Beyond scaling, dedicated feedback trained by strictly local rules provides a controlled design space for testing additional biologically motivated motifs (e.g., alternative reconstruction targets, partial or gated feedback, and sparsity constraints) without introducing global gradient transport.
In the present architecture, each feedback pathway is trained to reconstruct the activity of its paired forward layer, including reconstruction of the input representation as a special case. This explicit reconstruction signal suggests a practical role in settings that require online adaptation under locality constraints: feedback can carry short-horizon context and provide internal consistency signals while synaptic updates remain strictly local and do not require backpropagation through time. A longer-term extension is to couple this reconstruction pathway to a persistent storage mechanism, so that previously encountered internal states can be reinstated from partial cues across extended time—an episodic-memory-like operation. Importantly, the current model does not implement storage, temporal indexing, or capacity control; evaluating this direction will require sequence-level protocols (e.g., delayed recall and replay-like consolidation) and longer-horizon measurements of stability–plasticity trade-offs under sustained adaptation.
Another direction is hybridization with gradient-based training. A conservative use is to treat local feedback–reconstruction training as an initialization or auxiliary pretraining phase that shapes representations and stabilizes dynamics, followed by conventional gradient-based optimization for task-level performance. More broadly, the reconstruction pathway could also be leveraged as an auxiliary objective or regularizer for a separate learner, but the appropriate scheduling and the regimes in which hybrid training is beneficial remain open questions.